OpenStudy (anonymous):

Simplify the following expression: (5x2 + 3x + 4) - (2x2 - 6x + 3). If the final answer is written in the form Ax2 + Bx + C, what is the value of B?

OpenStudy (ivettef365):

first take out of parenthesis, so you have 5x^2+3x+4-2x^2+6x-3 now combine like terms

OpenStudy (anonymous):

so 5x,3x,2x, and 6x then you have 4,2, and 3 right?

OpenStudy (anonymous):

this algebra man

OpenStudy (ivettef365):

you can combine 5x^2 -2x^2 = 3x^2 then combine 3x + 6x = ?? and finally 4 - 3 = ??

OpenStudy (anonymous):

3x+6x=9x and 4-3=1 so you'd subtract those two yea? Im sorry im really bad at this

OpenStudy (ivettef365):

ok so you have as the answer 3x^2+9x+1

OpenStudy (ivettef365):

so what is the value of B

OpenStudy (anonymous):

is it 19?

OpenStudy (ivettef365):

no, the number that is in the B position is 9 , look at the formula and look at the answer

OpenStudy (anonymous):

So B is 9?

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

Ax^2 + Bx + C and you have 3x^2 + 9x +1

OpenStudy (ivettef365):

A = 3, B = 9 and C = 1

OpenStudy (ivettef365):

understand ?

OpenStudy (anonymous):

Yeah thank you so much

OpenStudy (anonymous):

How do i give you a medal?

OpenStudy (ivettef365):

just hit where it say "best response"

OpenStudy (anonymous):

Thank you

OpenStudy (ivettef365):

yw :)